A particle moves in circle of radius R=21/22m with const. speed of 1m/s. Find:


a) Magnitude of avge.velocity=?


b)Magnitude of avge.Acceleration in 2sec=?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

distance travell in 2 sec = 1* 2 =2m

   angel made in 2 sec = 2/21/22 radian

                                =2*22/21  * 180 *7/22  degree

                               = 120 degree

avg  velocity = total distance /time

                  = (√3 /2  * 21/22)/2  =.413  m/sec

avg acceleration  = change in velocity /time

                        =| 1 i  -(-√3 /2 i - 1/2 j)|/2

                        =|3/2 i+√3 /2j|/2

                        =√3/4  m/sec2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin